# Solving the Poisson equation using coupled Markov chains

### v3

@@ Abstract
This article shows how coupled Markov chains that meet exactly after a random number of iterations can be used to generate unbiased estimators of the solutions of the Poisson equation. We establish connections to recently-proposed unbiased estimators of Markov chain equilibrium expectations. We further employ the proposed estimators of solutions of the Poisson equation to construct unbiased estimators of the asymptotic variance of Markov chain ergodic averages, involving a random but finite computing time. The proposed estimators are studied under realistic assumptions on the meeting times of the coupled chains and on the existence of moments of test functions under the target distribution. We provide experiments in toy examples as well as more realistic settings.
@@ 1.1 Central Limit Theorem and the Poisson equation
Markov chain Monte Carlo (MCMC) methods form a convenient family of simulation techniques with many applications in statistics. With $(\mathbb{X},\mathcal{X})$ a measurable space and $\pi$ a probability measure of statistical interest, MCMC involves simulation of a discrete-time, time-homogeneous Markov chain $X=(X_{t})_{t\geq 0}$, with a $\pi$-invariant Markov transition kernel $P$ and some initial distribution $\pi_{0}$. Letting $L^{p}(\pi)=\{f:\pi(\left|f\right|^{p})<\infty\}$, where $\pi(f)=\int f(x)\pi({\rm d}x)$, the interest is to approximate $\pi(h)$ for some function $h\in L^{1}(\pi)$, termed the test function.
In particular, after simulating the chain until time $t$, one may approximate an integral of interest $\pi(h)$ via the ergodic average $t^{-1}\sum_{s=0}^{t-1}h(X_{s})$. Under weak assumptions, such averages converge almost surely to $\pi(h)$ as $t\to\infty$, and under stronger but still realistic assumptions on $\pi_{0}$, $P$ and $h$ (see Theorem 3), they satisfy central limit theorems (CLTs),
@@
where $v(P,h)$ is the asymptotic variance associated with the Markov kernel $P$ and the function $h$ (see, e.g., Jones, 2004). One standard route to proving a CLT is via a solution of the Poisson equation for $h$ associated with $P$, i.e. any function $g$ such that
@@
where $P$ is viewed as a Markov operator, i.e. $Pg:x\mapsto\int P(x,{\rm d}y)g(y)$ and $I$ is the identity. In particular, if $h,g\in L^{2}(\pi)$ then (1.1) holds with initial distribution $\pi$ by Douc et al. (2018, Theorem 21.2.5), and
@@
If $g_{\star}$ is well-defined, then it is straightforward to check that $(I-P)g_{\star}=h_{0},$ so that $g_{\star}$ is indeed a solution to the Poisson equation (1.2). For brevity, we will call solutions to (1.2) fishy functions. Fishy functions are not unique, since $g_{\star}+c$ is also fishy for any constant $c\in\mathbb{R}$. Moreover, if $g_{\star}\in L^{1}(\pi)$ (see Theorem 3) then $\pi(g_{\star})=0$, and by Douc et al. (Lemma 21.2.2, 2018) all fishy functions $g$ are equal to $g_{\star}$ up to an additive constant. Inserting the function $g_{\star}$ in place of $g$ in (1.3) leads to the more familiar expression
@@
where the subscript $\pi$ indicates that $X_{0}\sim\pi$. That expression can be obtained with direct calculations from $\lim_{t\to\infty}{\rm var}_{\pi}(t^{-1/2}\sum_{s=0}^{t-1}h(X_{s}))$.
@@ 1.2 Couplings of Markov chains
Our contributions rely on couplings of Markov chains. For any two distributions $\mu$ and $\nu$ on $(\mathbb{X},\mathcal{X})$, a coupling refers to a pair of random variables $(X,Y)$ on $(\mathbb{X}\times\mathbb{X},\mathcal{X}\times\mathcal{X})$ such that $X\sim\mu$ and $Y\sim\nu$. This extends to Markov chains, viewed as distributions on the path space $(\prod_{j=0}^{\infty}\mathbb{X},\prod_{j=0}^{\infty}\mathcal{X})$. We will consider only pairs of chains that start from different initial distributions but evolve with the same transition $P$, and we will focus on Markovian couplings: for two Markov chains $(X_{t})$ and $(Y_{t})$ we will consider a joint process on $(\prod_{j=0}^{\infty}\mathbb{X}\times\mathbb{X},\prod_{j=0}^{\infty}\mathcal{X}\times\mathcal{X})$ that is Markov with transition kernel $\bar{P}$, and such that its first and second coordinates are distributed as $(X_{t})$ and $(Y_{t})$ respectively. Couplings are often used in the convergence analysis of Markov chains (e.g. Jerrum, 1998), and they play a role in some MCMC methods, for example in perfect sampling (Propp & Wilson, 1996), convergence diagnostics (Johnson, 1996), variance reduction (Neal & Pinto, 2001), and unbiased estimation (Glynn & Rhee, 2014; Jacob et al., 2020).
@@ 1.3 Contributions
Our contributions are as follows. In Section 2 we elicit links between couplings of Markov chains and the Poisson equation (1.2). The connection leads to new estimators of pointwise evaluations of the solutions. We establish the lack of bias of these estimators and the finiteness of their moments under conditions on the function $h$ and on the coupled transition $\bar{P}$. In Section 3 we re-derive the unbiased estimators of $\pi(h)$ pioneered by Glynn & Rhee (2014), and the variants of Jacob et al. (2020). We refer to these estimators of $\pi(h)$ as unbiased MCMC. This novel derivation leads to improved theoretical results that provide simple conditions for the finiteness of their moments of any order, which will be used in subsequent sections. We also relate the efficiency of unbiased MCMC with the efficiency of ergodic average MCMC, which is the inverse of the asymptotic variance $v(P,h)$. In Section 4 we propose estimators of the asymptotic variance $v(P,h)$, combining the estimators of fishy functions from Section 2 with unbiased MCMC as in Section 3. These estimators can be used to compare the efficiencies of unbiased and ergodic average MCMC, and to compare MCMC algorithms to one another. We experiment with the proposed estimators in Section 5 and Section 6 concludes.
@@ 2.1 Coupled Markov chains and meeting time
We consider a time-homogeneous, discrete-time Markov chain $(X,Y)=(X_{t},Y_{t})_{t\geq 0}$ with Markov kernel $\bar{P}$, such that $X=(X_{t})_{t\geq 0}$ and $Y=(Y_{t})_{t\geq 0}$ are both time-homogeneous, discrete-time Markov chains with kernel $P$. We use subscripts to denote the distribution of $(X_{0},Y_{0})$. For example $\mathbb{P}_{x,y}$ is the law of $(X,Y)$ when $(X_{0},Y_{0})=(x,y)$, and $\mathbb{P}_{\nu}$ indicates $(X_{0},Y_{0})\sim\nu$. When only one chain is referenced, e.g. $X$, we may similarly write $\mathbb{P}_{x}$ and $\mathbb{P}_{\mu}$ to indicate $X_{0}=x$ and $X_{0}\sim\mu$, respectively.
@@
is almost surely finite. Furthermore, we impose that $X_{t}=Y_{t}$ with probability $1$ for all times $t\geq\tau$. Johnson (1998); Jacob et al. (2020) and others have shown how to construct such couplings for some realistic MCMC algorithms; Appendix A provides some reminders and pointers.
We now introduce the main assumption in this manuscript: we require the meeting time $\tau$ to have $\kappa$ finite moments, with $\kappa>1$, for two chains that would independently start from the stationary distribution $\pi$, i.e. $(X_{0},Y_{0})\sim\pi\otimes\pi$.
@@
Interpretation and verification of this assumption are discussed in Section 2.3. The assumption is enough to justify the existence of the function $g_{\star}$ in Definition 1, to justify that $g_{\star}$ is solution to (1.2) in $L_{0}^{1}(\pi)=\{f\in L^{1}(\pi):\pi(f)=0\}$, and to establish the CLT (1.1) for a class of test functions.
@@
The proof in Appendix B.2 relies heavily on the strategy of Douc et al. (2018, Section 21.4.1) but features $g_{\star}$ from Definition 1 more prominently, and uses the CLT condition from Maxwell & Woodroofe (2000) rather than Douc et al. (2018, Theorem 21.4.1). Note that the CLT does not require $g_{\star}\in L_{0}^{2}(\pi)$. As a complement to Theorem 3, Theorem 27 in Appendix B.1 states that, if $m>\kappa/(\kappa-1)$, then $g_{\star}\in L_{0}^{p}(\pi)$ for $p\geq 1$ such that $p^{-1}>m^{-1}+\kappa^{-1}$.
@@ 2.2 Unbiased approximation of fishy functions
Due to their convenience for the theoretical analysis of Markov chains, solutions of the Poisson equation have been studied extensively (see, e.g., Glynn & Meyn, 1996; Glynn & Infanger, 2022). However, the Poisson equation is not analytically solvable for most Markov chains and functions of interest, and consistent approximations have been lacking. We next define a family of fishy functions that lends itself to estimation with coupled chains.
@@
For $y\in\mathbb{X}$, consider the function
@@
For $X_{0}=x\in\mathbb{X}$ and $Y_{0}=y$, the proposed estimator of $g_{y}(x)$ is
@@
where $(X,Y)$ and $\tau$ are defined in Section 2.1. In particular, the dependence of $G_{y}(x)$ in (2.4) on $x$ and $y$ is that the law of $(X,Y)$ is $\mathbb{P}_{x,y}$, and so we may omit the subscript $(x,y)$ when there is no ambiguity. We will also sometimes denote $G_{y}$ by $G$.
@@
since $h(X_{t})=h(Y_{t})$ for all $t\geq\tau$, and each term $h(X_{t})-h(Y_{t})$ has expectation $P^{t}h(x)-P^{t}h(y)$ under $\mathbb{P}_{x,y}$. Hence, $G_{y}(x)$ has expectation equal to $g_{y}(x)$ if we can justify the interchange of expectation and infinite sum. The following result, established in Appendix B.3 as a special case of Theorem 37, provides conditions for $G_{y}(x)$ to be unbiased for $g_{y}(x)$ and to have finite moments.
@@
For our subsequent asymptotic variance estimators in Section 4, it is important that this is a pointwise result. The random variable $G_{y}(x)$ may be simulated using Algorithm 2.1 with $L=0$. The cost of sampling $G_{y}(x)$ is the cost of running a pair of chains until they meet, which is typically comparable to twice the cost of running one chain for the same number of steps, i.e. $2\tau$.
@@
Output: coupled chains and the meeting time defined as the smallest $t$ at which $X_{t}=Y_{t-L}$.
@@ 2.3 Interpretation and verification of Assumption 2
Since Assumption 2 is used extensively in this manuscript, this section provides details on how to interpret and to verify it.
The assumption implies a polynomially decaying survival function $\mathbb{P}_{x,y}(\tau>t)$ of order $\kappa$ and, conversely, one may verify the assumption by showing that $\mathbb{P}_{x,y}(\tau>t)$ decays polynomially with order $s>\kappa$ with a dependence on $(x,y)$ that is not too strong. The proof of the next result is in Appendix B.1.
@@
If Assumption 2 holds then for all $t\geq 0$,
$\mathbb{P}_{x,y}(\tau>t)\leq\mathbb{E}_{x,y}[\tau^{\kappa}](t+1)^{-\kappa},$
and $\mathbb{E}_{x,y}[\tau^{\kappa}]<\infty$ for $\pi\otimes\pi$-almost all $(x,y)$. Conversely, if for some $s>\kappa$, there exists $\tilde{C}:\mathbb{X}\times\mathbb{X}\to\mathbb{R}$ with $\pi\otimes\pi(\tilde{C})<\infty$ such that for $\pi\otimes\pi$-almost all $(x,y)$, we have for all $t\geq 0$,
$\mathbb{P}_{x,y}(\tau>t)\leq\tilde{C}(x,y)(t+1)^{-s},$
@@
We can adapt results from Jacob et al. (2020) to verify Assumption 2 in the case of geometrically ergodic Markov chains. Assumption 9 below, referred to as a geometric drift condition, has been shown to hold for many MCMC algorithms, under various conditions on the target distribution, see for example Jarner & Hansen (2000) on Metropolis–Rosenbluth–Teller–Hastings (MRTH) with random walk proposals, Roberts & Tweedie (1996) on Langevin algorithms, and Durmus et al. (2020) on Hamiltonian Monte Carlo.
@@
The Markov kernel $P$ is $\pi$-invariant, $\varphi$-irreducible and aperiodic, and there exists a measurable function $V:\;\mathcal{X}\to[1,\infty)$, $\lambda\in(0,1)$, $b<\infty$ and a small set $\mathcal{C}$ such that for all $x\in\mathcal{X}$,
$PV(x)\leq\lambda V(x)+b\mathds{1}(x\in\mathcal{C}).$
While Assumption 9 is on the marginal transition $P$, the following assumption is on the coupling $\bar{P}$. It is for example satisfied for couplings of MRTH with Normal proposals, when the proposals are maximally coupled and when the set $\mathcal{C}$ is any compact set.
@@
There exist $\mathcal{C}$ and $\epsilon\in(0,1)$ such that
$\inf_{(x,y)\in\mathcal{C}\times\mathcal{C}}\bar{P}((x,y),\{x=y\})\geq\epsilon.$ (2.5)
The following result relates these assumptions to Assumption 2.
@@
Suppose that $P$ satisfies Assumption 9 with a small set $\mathcal{C}$ of the form $\mathcal{C}=\{x:V(x)\leq\bar{v}\}$, for which Assumption 10 holds with some $\epsilon>0$, and assume that $\lambda+2b/(1+\bar{v})<1$. Then Assumption 2 holds for all $\kappa>1$.
The above result is very similar to Proposition 4 of Jacob et al. (2020), and its proof is omitted. The only difference comes from the initialization from $\pi\otimes\pi$ rather than $\pi_{0}P\otimes\pi_{0}$; the explicit assumption $\pi_{0}(V)<\infty$ in Proposition 4 of Jacob et al. (2020) becomes $\pi(V)<\infty$, but this always holds under Assumption 9, as stated in Lemma 14.1.10 of Douc et al. (2018). Proposition 4 of Jacob et al. (2020) was used to show that the meeting time had Geometric tails for couplings of Hamiltonian Monte Carlo in Heng & Jacob (2019); Xu et al. (2021), and for couplings of the Bouncy Particle Sampler in Corenflos et al. (2023), under assumptions similar to strong log-concavity and smoothness of the target distribution. It was also used in Biswas et al. (2022) for couplings of a Gibbs sampler for high-dimensional regression with shrinkage priors, which we employ in Section 5.3. The tails of the meeting times can also be studied directly, on a case-by-case basis, as is done in Jacob et al. (2019); Lee et al. (2020) for couplings of conditional particle filters, in Nguyen et al. (2022) for couplings of a Gibbs sampler for Bayesian clustering with Dirichlet process mixtures, and in Section 2.4 below for a coupling of a Gibbs sampler in a Cauchy location model.
The reasoning underlying Proposition 11 can be generalized to polynomially ergodic chains, for which Assumption 2 holds for only some values of $\kappa$; see Theorem 2 in Middleton et al. (2020). We provide a variant of this result below, with the proof in Appendix B.1.
@@
The Markov kernel $P$ is $\pi$-invariant and there exist a measurable function $V:\;\mathcal{X}\to[1,\infty)$, $\alpha\in(0,1)$, $\vartheta,b<\infty$ and a small set $\mathcal{C}$ such that we have $\sup_{\mathcal{C}}V<\infty$ and $\inf_{{\mathcal{C}}^{\complement}}V^{\alpha}>b/\vartheta$ and for all $x\in\mathcal{X}$,
$PV(x)\leq V(x)-\vartheta V^{\alpha}(x)+b\mathds{1}(x\in\mathcal{C}).$

### v4

@@ Abstract
This article shows how coupled Markov chains that meet exactly after a random number of iterations can be used to generate unbiased estimators of the solutions of the Poisson equation. Through this connection, we re-derive known unbiased estimators of expectations with respect to the stationary distribution of a Markov chain and provide conditions for the finiteness of their moments. We further construct unbiased estimators of the asymptotic variance of Markov chain ergodic averages, and provide conditions for the finiteness of the estimators’ moments of any order. If their second moment is finite, the average of independent copies of such estimators converges to the asymptotic variance at the Monte Carlo rate, comparing favorably to known rates for batch means and spectral variance estimators. The results are illustrated with numerical experiments.
@@ 1.1 Central Limit Theorem and the Poisson equation
Markov chain Monte Carlo (MCMC) methods form a convenient family of simulation techniques with many applications in statistics. With $(\mathbb{X},\mathcal{X})$ a measurable space and $\pi$ a probability measure of statistical interest, MCMC involves simulation of a discrete-time, time-homogeneous Markov chain $X=(X_{t})_{t\geq 0}$, with a $\pi$-invariant Markov transition kernel $P$ and initial distribution $\pi_{0}$. Letting $L^{p}(\pi)=\{f:\pi(\left|f\right|^{p})<\infty\}$, where $\pi(f)=\int f(x)\pi({\rm d}x)$, the interest is to approximate $\pi(h)$ for some function $h\in L^{1}(\pi)$, termed the test function.
In particular, after simulating the chain until time $t$, one may approximate an integral of interest $\pi(h)$ via the ergodic average $t^{-1}\sum_{s=0}^{t-1}h(X_{s})$. Under weak assumptions, such averages converge almost surely to $\pi(h)$ as $t\to\infty$ (see for example Theorem 17.0.1 in Meyn & Tweedie, 2009), and under stronger but still realistic assumptions on $\pi_{0}$, $P$ and $h$, ergodic averages satisfy central limit theorems (CLTs),
@@
where $v(P,h)$ is the asymptotic variance associated with the Markov kernel $P$ and the function $h$ (see for example Theorem 1 in Jones, 2004). One standard route to proving a CLT is via a solution of the Poisson equation for $h$ associated with $P$, i.e. any function $g$ such that
@@
where $P$ is viewed as a Markov operator, i.e. $Pg:x\mapsto\int P(x,{\rm d}y)g(y)$ and $I$ is the identity. Theorem 3 below (or, for example, Theorem 21.2.5 in Douc et al., 2018) provides conditions for (1.1) to hold and for the asymptotic variance to be of the form
@@
If $g_{\star}$ is well-defined, then it is straightforward to check that $(I-P)g_{\star}=h_{0},$ so that $g_{\star}$ is indeed a solution to the Poisson equation (1.2). For brevity, we will call solutions to (1.2) fishy functions. Fishy functions are not unique, since $g_{\star}+c$ is also fishy for any constant $c\in\mathbb{R}$. Moreover, if $\pi$ is the unique invariant distribution of $P$ and if $g_{\star}\in L^{1}(\pi)$ (see Theorem 3) then $\pi(g_{\star})=0$, and by Douc et al. (Lemma 21.2.2, 2018) all fishy functions $g$ are equal to $g_{\star}$ up to an additive constant. Inserting the function $g_{\star}$ in place of $g$ in (1.3) leads to the more familiar expression
@@
where the subscript $\pi$ indicates that $X_{0}\sim\pi$. That expression is retrieved with direct calculations from developing ${\rm var}_{\pi}(t^{-1/2}\sum_{s=0}^{t-1}h(X_{s}))$ and taking the limit as $t\to\infty$.
@@ 1.2 Couplings of Markov chains
Our contributions rely on couplings of Markov chains. For any two distributions $\mu$ and $\nu$ on $(\mathbb{X},\mathcal{X})$, a coupling refers to a pair of random variables $(U,V)$ on $(\mathbb{X}\times\mathbb{X},\mathcal{X}\otimes\mathcal{X})$ such that $U\sim\mu$ and $V\sim\nu$. This extends to Markov chains, viewed as distributions on the path space $(\prod_{t=0}^{\infty}\mathbb{X},\bigotimes_{t=0}^{\infty}\mathcal{X})$. We consider pairs of chains that start from different distributions but evolve with the same transition $P$, and we will focus on Markovian couplings: for two Markov chains $(X_{t})_{t\geq 0}$ and $(Y_{t})_{t\geq 0}$ we will consider a joint process on $(\prod_{t=0}^{\infty}\mathbb{X}\times\mathbb{X},\bigotimes_{t=0}^{\infty}\mathcal{X}\otimes\mathcal{X})$ that is Markov with transition kernel $\bar{P}$, and such that its first and second coordinates are distributed as $(X_{t})_{t\geq 0}$ and $(Y_{t})_{t\geq 0}$ respectively.
@@ 1.3 Contributions
In Section 2 we elicit links between couplings of Markov chains and the Poisson equation (1.2). The connection leads to estimators of pointwise evaluations of the solutions. We establish the lack of bias of these estimators and the finiteness of their moments under conditions on the function $h$ and on the coupled transition $\bar{P}$. In Section 3 we re-derive the unbiased estimators of $\pi(h)$ pioneered by Glynn & Rhee (2014), and the variants of Jacob et al. (2020). We refer to these estimators of $\pi(h)$ as unbiased MCMC. Our derivation leads to simple conditions for the finiteness of their moments of any order (as opposed to second moments only in the aforementioned articles), which will be used in subsequent sections. We also relate the efficiency of unbiased MCMC with the efficiency of ergodic average MCMC, which is the inverse of the asymptotic variance $v(P,h)$. In Section 4 we propose estimators of the asymptotic variance $v(P,h)$, combining the estimators of fishy functions from Section 2 with unbiased MCMC as in Section 3. The proposed estimators of $v(P,h)$ are built from independent runs of coupled chains of random lengths, instead of long runs. They converge to $v(P,h)$ at the Monte Carlo rate, which is faster than known rates of convergence for batch means and spectral variance estimators. Estimators of $v(P,h)$ can be used to compare the efficiencies of unbiased and ergodic average MCMC, and to compare MCMC algorithms to one another without ever relying on long run asymptotics. We experiment with the proposed estimators in Section 5 and Section 6 concludes.
@@ 2.1 Coupled Markov chains and meeting time
We consider a time-homogeneous, discrete-time Markov chain $(X,Y)=(X_{t},Y_{t})_{t\geq 0}$ with Markov kernel $\bar{P}$, such that $X=(X_{t})_{t\geq 0}$ and $Y=(Y_{t})_{t\geq 0}$ are both time-homogeneous, discrete-time Markov chains with kernel $P$. We use subscripts to denote the distribution of $(X_{0},Y_{0})$. For example $\mathbb{P}_{x,y}$ is the law of $(X,Y)$ when $(X_{0},Y_{0})=(x,y)$, and $\mathbb{P}_{\bar{\nu}}$ indicates $(X_{0},Y_{0})\sim\bar{\nu}$, for some distribution $\bar{\nu}$ on the joint space. When only one chain is referenced, e.g. $X$, we may similarly write $\mathbb{P}_{x}$ and $\mathbb{P}_{\mu}$ to indicate $X_{0}=x$ and $X_{0}\sim\mu$, respectively.
@@
is almost surely finite. Furthermore, we impose that $X_{t}=Y_{t}$ with probability $1$ for all $t\geq\tau$. Johnson (1998); Jacob et al. (2020) and others have shown how to construct such couplings for some realistic MCMC algorithms; Appendix A.1 provides pointers.
We now introduce the main assumption in this manuscript: we require the meeting time $\tau$ to have $\kappa$ finite moments, with $\kappa>1$, for two chains that would independently start from the stationary distribution $\pi$, i.e. $(X_{0},Y_{0})\sim\pi\otimes\pi$, the independent coupling of $\pi$ with itself.
@@
Interpretation and verification of this assumption are discussed in Section 2.3. The assumption that $P$ is $\pi$-irreducible is equivalent to assuming that it has an irreducibility measure, since an invariant probability measure is a maximal irreducibility measure (Douc et al., 2018, Theorem 9.2.15) and implies that $\pi$ is necessarily the unique invariant probability measure for $P$ (Douc et al., 2018, Corollary 9.2.16). Assumption 2 is sufficient to guarantee that the chain is aperiodic (Appendix B.1), to justify the existence of the function $g_{\star}$ in Definition 1, to justify that $g_{\star}$ is solution to (1.2) in $L_{0}^{1}(\pi)=\{f\in L^{1}(\pi):\pi(f)=0\}$, and to establish the CLT (1.1) for a class of test functions.
@@
The proof in Appendix B.2 relies heavily on the strategy of Douc et al. (2018, Section 21.4.1) but features $g_{\star}$ from Definition 1 more prominently, and uses the CLT condition from Maxwell & Woodroofe (2000) rather than Douc et al. (2018, Theorem 21.4.1). Note that the CLT does not require $g_{\star}\in L_{0}^{2}(\pi)$. As a complement to Theorem 3, Theorem B.4 states that, if $m>\kappa/(\kappa-1)$, then $g_{\star}\in L_{0}^{p}(\pi)$ for $p\geq 1$ such that $p^{-1}>m^{-1}+\kappa^{-1}$.
@@ 2.2 Unbiased approximation of fishy functions
Solutions of the Poisson equation have been studied extensively (see, e.g., Glynn & Meyn, 1996; Glynn & Infanger, 2022). However, the Poisson equation is not analytically solvable for most Markov chains and functions of interest, and consistent approximations have been lacking. We consider a family of fishy functions that are amenable to estimation using coupled chains.
@@
For $y\in\mathbb{X}$ and $g_{\star}$ in Definition 1, define the function
@@
For $x,y\in\mathbb{X}$, the proposed estimator of $g_{y}(x)$ is
@@
where $(X,Y)$ is a Markov chain starting from $(X_{0},Y_{0})=(x,y)$, evolving according to $\bar{P}$ defined in Section 2.1, and $\tau$ is defined in (2.2). The random variable $G_{y}(x)$ is thus a function of the coupled chains $(X,Y)$ from time zero to $\tau-1$. We will sometimes denote $G_{y}$ by $G$.
@@
since $h(X_{t})=h(Y_{t})$ for all $t\geq\tau$, and each term $h(X_{t})-h(Y_{t})$ has expectation $P^{t}h(x)-P^{t}h(y)$ under $\mathbb{P}_{x,y}$. Hence, $\mathbb{E}[G_{y}(x)]$ is equal to $g_{y}(x)$ if we can justify the interchange of expectation and infinite sum. The following result, established in Appendix B.3 provides conditions for $G_{y}(x)$ to be unbiased and to have finite moments.
@@
For our subsequent asymptotic variance estimators in Section 4, it is important that this result is pointwise. The random variable $G_{y}(x)$ may be simulated using Algorithm 2.1 with $L=0$ (setting $L\geq 1$ will be useful from Section 3 onwards). The cost of sampling $G_{y}(x)$ is the cost of running a pair of chains until they meet, which is typically comparable to twice the cost of running one chain for the same number of steps, i.e. $2\tau$.
@@
Output: coupled chains and meeting time $\tau=\inf\{t>L:X_{t}=Y_{t-L}\}$.
@@ 2.3 Interpretation and verification of Assumption 2
Since we use Assumption 2 extensively, this section provides details on how to interpret and to verify it via drift conditions. The assumption implies a polynomially decaying survival function $\mathbb{P}_{x,y}(\tau>t)$ of order $\kappa$ and, conversely, one may verify the assumption by showing that $\mathbb{P}_{x,y}(\tau>t)$ decays polynomially with order $s>\kappa$ with a dependence on $(x,y)$ that is not too strong. The proof of the next result is in Appendix B.1.
@@
If Assumption 2 holds then
$\forall t\geq 0\quad\mathbb{P}_{x,y}(\tau>t)\leq\mathbb{E}_{x,y}[\tau^{\kappa}](t+1)^{-\kappa},$
and $\mathbb{E}_{x,y}[\tau^{\kappa}]<\infty$ for $\pi\otimes\pi$-almost all $(x,y)$. Conversely, if for some $s>\kappa$, there exists $\tilde{C}:\mathbb{X}\times\mathbb{X}\to\mathbb{R}$ with $\pi\otimes\pi(\tilde{C})<\infty$ such that for $\pi\otimes\pi$-almost all $(x,y)$, we have
$\forall t\geq 0\quad\mathbb{P}_{x,y}(\tau>t)\leq\tilde{C}(x,y)(t+1)^{-s},$
@@
We can adapt results from Jacob et al. (2020) to verify Assumption 2 in the case of geometrically ergodic Markov chains. Variants of the following assumption are referred to as a geometric drift condition. Assumption 9 is very similar to Condition $D_{\text{g}}(V,\lambda,b,C)$, Definition 14.1.5 in Douc et al. (2018). Such conditions have been shown to hold for many combinations of MCMC algorithms and target distributions, see for example Jarner & Hansen (2000) on Metropolis–Rosenbluth–Teller–Hastings (MRTH) with random walk proposals, Roberts & Tweedie (1996) on Langevin algorithms, and Durmus et al. (2020) on Hamiltonian Monte Carlo.
@@
The Markov kernel $P$ is $\pi$-invariant, $\pi$-irreducible, and there exists a measurable function $V:\;\mathbb{X}\to[1,\infty)$, $\lambda\in(0,1)$, $b\in(0,\infty)$ and a small set $\mathcal{C}$ such that
$\forall x\in\mathbb{X}\quad PV(x)\leq\lambda V(x)+b\mathds{1}(x\in\mathcal{C}).$
While Assumption 9 is on the transition $P$, the next assumption is on the coupling $\bar{P}$.
@@
There exist $\mathcal{C}$ and $\epsilon\in(0,1)$ such that, with $\mathcal{D}=\{(x,y)\in\mathbb{X}\times\mathbb{X}:x=y\}$,
$\inf_{(x,y)\in\mathcal{C}\times\mathcal{C}}\bar{P}((x,y),\mathcal{D})\geq\epsilon.$ (2.5)
Assumption 10 states that meeting occurs in one step with probability at least $\epsilon$ when both chains are simultaneously in $\mathcal{C}$ and evolve according to $\bar{P}$. It is for example satisfied for couplings of Metropolis–Rosenbluth–Teller–Hastings (MRTH) with Normal proposals using the reflection-maximal coupling described in Appendix A.1 for any bounded set $\mathcal{C}$ on which the target density is upper bounded.
@@
Suppose that $P$ satisfies Assumption 9 with a small set $\mathcal{C}$ of the form $\mathcal{C}=\{x:V(x)\leq\bar{v}\}$, for some $\bar{v}\geq 1$, for which Assumption 10 holds with some $\epsilon>0$, and assume that $\lambda+2b/(1+\bar{v})<1$. Then Assumption 2 holds for all $\kappa>1$.
The above result is very similar to Proposition 4 of Jacob et al. (2020), and its proof is omitted. The only difference comes from the initialization from $\pi\otimes\pi$ rather than $\pi_{0}P\otimes\pi_{0}$, where $\pi_{0}$ is the initial distribution of the chain. The explicit assumption $\pi_{0}(V)<\infty$ in Proposition 4 of Jacob et al. (2020) becomes $\pi(V)<\infty$, but this always holds under Assumption 9, as stated in Lemma 14.1.10 of Douc et al. (2018). As noted in Jacob et al. (2020), if sub-level sets $\mathcal{C}(v)=\{x:V(x)\leq v\}$ are small for all sufficiently large $v$ and Assumption 9 holds with $\mathcal{C}=\mathcal{C}(v)$, then Assumption 9 also holds for $\mathcal{C}=\mathcal{C}(\bar{v})$ with $\bar{v}\geq v$ and so the condition $\lambda+2b/(1+\bar{v})<1$ can be satisfied by taking $\bar{v}$ large enough, while the quantity $\epsilon$ in Assumption 10 typically decreases but remains positive.
Proposition 4 of Jacob et al. (2020) was used to show that the meeting time had geometric tails for couplings of Hamiltonian Monte Carlo in Heng & Jacob (2019, Theorem 1), and for couplings of the Bouncy Particle Sampler in Corenflos et al. (2023, Proposition 3.1), under assumptions similar to strong log-concavity and smoothness of the target. It was also used in Biswas et al. (2022, Proposition 2) for couplings of a Gibbs sampler for high-dimensional regression with shrinkage priors. Thus Proposition 11 can similarly be applied to these settings where geometric drift conditions have been shown to hold.
To cover other cases, the reasoning can be extended to polynomially ergodic chains, for which Assumption 2 may hold only for some values of $\kappa$. We consider the following assumption, which is a variant of a polynomial drift condition. The assumption is very similar to Condition $D_{\text{sg}}(V,\phi,b,C)$, Definition 16.1.7 in Douc et al. (2018), but includes the condition $\inf_{{\mathcal{C}}^{\complement}}V^{\alpha}>b/\vartheta$. As discussed above, this condition can be met if sub-level sets are small.
@@
The Markov kernel $P$ is $\pi$-invariant, $\pi$-irreducible, and there exists a measurable function $V:\;\mathbb{X}\to[1,\infty)$, $\alpha\in(0,1)$, $\vartheta,b\in(0,\infty)$ and a small set $\mathcal{C}$ such that we have $\sup_{\mathcal{C}}V<\infty$ and $\inf_{{\mathcal{C}}^{\complement}}V^{\alpha}>b/\vartheta$ and
$\forall x\in\mathbb{X}\quad PV(x)\leq V(x)-\vartheta V^{\alpha}(x)+b\mathds{1}(x\in\mathcal{C}).$
Under Assumptions 10-12 we obtain the next result, established in Appendix B.1.